Using the formula for gravity, find the force of gravity on a 0.70-kg mass at Earth's surface. (The mass of Earth is 6×1024kg, and its radius is 6.4×106m.).
1
Expert's answer
2019-09-27T09:43:21-0400

The Newton's law of universal gravitation states


"F=G\frac{mM}{r^2}"

So


"F=6.67\times 10^{-11}\:\rm N\cdot m^2/kg^2\times \frac{0.70\:\rm kg\times 6\times 10^{24}\: kg}{(6.4\:\times10^6 \:\rm{m})^2}"

"=6.84\:\rm N"